Calculate the unweighted average and median for the following companies' wages. Also calculate the weighted average for both companies' wages together.

Company A - $8.00, $8.15, $8.25, $9.00, $8.00

Company B - $8.25, $10.00, $8.75, $10.25, $8.45, $8.50, $8.00
Answer #1

Answer:

Median:

  • Company A wages in increasing order :
    • 8.00 , 8.00 , 8.15 , 8.25 , 9.00
    • So the median wage for company A = Central value = $8.15     (Answer)
  • Company B wages in increasing order :
    • 8.00 , 8.25 , 8.45 , 8.50 , 8.75, 10.00 , 10.25
    • So the median wage for company B = Central value = $8.50     (Answer)

Average:

  • Company A average salary = (8.00 + 8.00 + 8.15 + 8.25 + 9.00) / 5 = $ 8.28      (Answer)
  • Company B average salary = (8.00 + 8.25 + 8.45+8.50 + 8.75 + 10.00 +10.25 ) / 7 = $8.8857

Weighted average:

  • Considering equal weightage for both the companies A&B.
    • Weighted average wages = (0.5* average of A) + (0.5* average of B)
    • Weighted Average = (0.5* 8.28) + (0.5 * 8.8857) = $ 8.5828 (Answer)